# A Theoretical Framework for Iteratively Enhanced Room-Temperature Single-Photon Detection

### v3

@@ III.1.4 Auxiliary Sources
Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, we note that implementing the ESPD scheme may not require single photon sources. The key parameter $P$ represents the detection probability of the state $|1\rangle$ from the auxiliary path with an ideal SPD when the control signal input is non-vacuum, regardless of the exact number of photons or the presence of unwanted $|0\rangle$. Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more opportunities for detecting the required $|1\rangle$ [^3] Since the employment of the filter block, for a non-vacuum input of a controlled module, if the auxiliary multi-photon signal results in more state $|1\rangle$, then a positive report has a higher chance to be produced, which provides benefits since it enlarges $P$, while, if the auxiliary multi-photon signal adds state $|0\rangle$, then they will be filtered out, thus providing no effects. Similarly, for a vacuum input of a controlled module, only $|0\rangle$ is added in auxiliary sources, which will be filtered out and thereby have no effect even if the source is multiple-photon.. However, the formulation of this, including its impact on controlled operations, needs a rigorous discussion and is reserved for future work.
III.2 Limitations and Future Works
@@
III.2.1 Theoretical Completeness
The theoretical framework established in Section II.2 remains an area for deeper exploration. While the numerical designs presented in Section II.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest. Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required. Given the inherent mathematical difficulty of such a non-linear and time-variant system and the sufficiency for practical technology employments by numerical designs, these specialized mathematical studies are reserved for the future.
III.2.2 Resource Overhead vs. Environmental Simplicity
Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section II.4
@@ III.2.1 Theoretical Completeness
While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels. Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections. Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD. For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (if not applying techniques such as multiplexing), resulting in hundreds of integrated components. While not posing a fundamental physical barrier (as discussed in Section III.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
However, this complexity is well within the capabilities of modern large-scale photonic integrated circuits. Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57]. Notably, very-large-scale integration of quantum photonics has been experimentally realized with over 2,500 components on a single silicon-on-insulator chip [58], verifying that the component density required for a multi-level ESPD is manufacturable using standard complementary metal-oxide-semiconductor processes. Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss material platforms. For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
III.2.3 Optimization Design
The inherent exponential resource scaling discussed in Section III.2.2 necessitates optimization for future practical deployments. Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$). Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence. As shown in Section II.4, the optimal choice of parameters $n$ (number of auxiliary signals in each level) and $k$ (decision threshold) is crucial: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost. Furthermore, as shown in Table 7 and visualized in Fig.7, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
@@ III.2.3 Optimization Design
Figure 7: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different $k$. Para 3 is the same as in Fig.2.
@@
The optimization problem also encompasses the optimal circuit topology. To mitigate cumulative error propagation, minimizing the maximal circuit length is essential. As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the paradigm: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section II.1. However, these architectural changes introduce new theoretical challenges: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
III.2.4 Physical Integrations and Experiments
@@
Despite the convincing indication of effectiveness via numerical simulations in Section II.4, physical implementations of the ESPD scheme require dedicated experimental validations. Reassuringly, as established in Section III.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization. However, the realization of the large-scale iterative circuit demands significant and focused engineering efforts, particularly concerning the precise integration and synchronization of the cascaded components. Furthermore, the seamless integration of the enhanced detector into various quantum optical tasks (such as QKD) needs full investigation. Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
III.2.5 Count Rate and Timing Jitter
@@
While the ESPD framework significantly enhances DE and suppresses DCR, it generally inherits the timing characteristics of the baseline detectors. Specifically, the system’s timing jitter and maximum count rate remain constrained by the intrinsic performance of the initial SPDs (e.g., the avalanche buildup time and dead time of InGaAs/InP devices). However, the architecture does not impose significant additional penalties: the optical-controlled quantum operations function on timescales usually faster than the electronic response of room-temperature SPDs, ensuring negligible added temporal broadening. Furthermore, as the final SPDs operate in parallel, the effective system throughput remains comparable to that of the constituent devices. Nevertheless, future iterations of the ESPD scheme could incorporate techniques such as multiplexing strategies to optimize these temporal metrics.
IV Conclusion
We have introduced the enhanced single-photon detector (ESPD) framework, which reformulates single-photon detection as an iteratively enhanced process based on integrating state preparations, controlled operations, projective measurements, and multi-copy analysis, rather than as a fixed device-level operation. Within this system-level theoretical framework, detector performance is progressively improved through structured quantum processing using exclusively non-superconducting, room-temperature components. Numerical simulations indicate that, starting from a conventional non-superconducting SPD, the ESPD framework can substantially enhance DE while strongly suppressing DCR, yielding effective performance metrics comparable to those of state-of-the-art superconducting SPDs, and significantly relaxing the minimal tolerable channel transmission rate in quantum communication. The ESPD framework makes explicit the trade-off between resource overhead and operating conditions, replacing cryogenic infrastructure with circuit integration complexity. While the experimental realization of large-scale ESPD architectures will require substantial engineering effort, the parameter studies presented in this work, grounded in the performance of existing photonic components, support the in-principle viability of the framework. In essence, this work establishes a system-level perspective on photon detection, showing that intrinsic device limitations can be transcended through iterative quantum processing and providing a principled foundation for the development of high-performance, room-temperature SPDs.
Data and Code Availability
@@ IV Conclusion
Author Contributions
Hao Shu (the unique author): study design, investigation, experiment, analysis, manuscript preparation.
Declarations
The author declares no competing interests.

### v4

@@ 3.1.4 Auxiliary Sources
Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, implementing the ESPD scheme may not require single photon sources. The key parameter $P$ represents the detection probability of the state $|1\rangle$ from the auxiliary path with an ideal SPD when the control signal input is non-vacuum, regardless of the exact number of photons or the presence of unwanted $|0\rangle$. Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more opportunities for detecting the required $|1\rangle$ [^3] Since the employment of the filter block, for a non-vacuum input of a controlled module, if the auxiliary multi-photon signal results in more state $|1\rangle$, then a positive report has a higher chance to be produced, which provides benefits since it enlarges $P$, while, if the auxiliary multi-photon signal adds state $|0\rangle$, then they will be filtered out, thus providing no effects. Similarly, for a vacuum input of a controlled module, only $|0\rangle$ is added in auxiliary paths, which will be filtered out and thereby have no effect even if the source is multiple-photon.. However, a rigorous formulation of this effect, including its impact on the controlled operations, remains an open problem and is left for future investigation.
3.2 Limitations and Future Works
@@
3.2.1 Theoretical Completeness
The theoretical framework established in Section 2.2 remains an area for deeper exploration. While the numerical designs presented in Section 2.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest. Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required. Given the inherent mathematical difficulty of this non-linear and time-variant system, and the fact that numerical designs already suffice for practical implementations, such specialized theoretical investigations are deferred to future work.
3.2.2 Resource Overhead vs. Environmental Simplicity
While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels. Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections. Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD. For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (in the absence of techniques such as multiplexing), resulting in hundreds of integrated components. While not posing a fundamental physical barrier (as discussed in Section 3.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
However, this complexity is well within the capabilities of modern large-scale photonic integrated circuits. Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57]. Notably, very-large-scale integration of quantum photonic components has already been experimentally demonstrated, with over 2,500 integrated on a single silicon-on-insulator chip [58], verifying that the component density required for a multi-level ESPD is manufacturable using standard complementary metal-oxide-semiconductor processes. Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss material platforms. For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
3.2.3 Optimization Design
Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section 2.4
@@ 3.2.3 Optimization Design
Figure 5: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different $k$. Para 3 is the same as in Table.2.
The inherent exponential resource scaling discussed in Section 3.2.2 necessitates optimization for future practical deployments. Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$). Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence. As demonstrated in Section 2.4, the choice of design parameters, namely $n$ (the number of auxiliary signals per level) and $k$ (the decision threshold), plays a critical role: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost. Furthermore, as shown in Table 7 and visualized in Fig.5, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
@@
The optimization problem also encompasses the optimal circuit topology. To mitigate cumulative error propagation, minimizing the maximal circuit length is essential. As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the paradigm: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section 2.1. However, these architectural changes introduce new theoretical challenges: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
3.2.4 Physical Integrations and Experiments
@@
Despite the convincing evidence provided by the numerical simulations in Section 2.4, the physical implementation of the ESPD scheme still requires dedicated experimental validation. Reassuringly, as established in Section 3.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization. However, the realization of the large-scale iterative circuit demands significant and focused engineering efforts, particularly concerning the precise integration and synchronization of the cascaded components. Furthermore, the seamless integration of the enhanced detector into various quantum optical tasks (such as QKD) needs full investigation. Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
3.2.5 Count Rate and Timing Jitter
@@
While the ESPD framework significantly enhances DE and suppresses DCR, it generally inherits the timing characteristics of the baseline detectors. Specifically, the system’s timing jitter and maximum count rate remain constrained by the intrinsic performance of the initial SPDs (e.g., the avalanche buildup time and dead time of InGaAs/InP devices). However, the ESPD architecture does not introduce significant additional penalties in principle: the optical-controlled quantum operations function on timescales usually faster than the electronic response of room-temperature SPDs, ensuring negligible added temporal broadening. Furthermore, as the final SPDs operate in parallel, the effective system throughput remains comparable to that of the constituent devices. Nevertheless, future iterations of the ESPD scheme could incorporate techniques such as multiplexing strategies to optimize these temporal metrics.
4 Conclusion
In summary, the enhanced single-photon detection (ESPD) framework is established to reformulate photon detection as an iteratively enhanced quantum-information-processing task composed of state preparation, controlled operations, projective measurements, and multi-copy decision analysis, rather than as a fixed device-level operation. Within this system-level theoretical framework, detection performance is progressively improved through structured quantum processing using exclusively non-superconducting, room-temperature components.
Numerical simulations, grounded in experimentally reported parameters and commercially available photonic components, demonstrate that the ESPD framework can substantially enhance the effective detection efficiency (DE) (from 59% to 93%) while strongly suppressing the effective dark count rate (DCR) (from $10^{-2}$ to $10^{-9}$) even when initialized with a conventional room-temperature SPD reported two decades ago. The resulting effective performance metrics are comparable to those of state-of-the-art superconducting single-photon detectors, significantly relaxing the minimal tolerable channel transmission rate in quantum communication applications.
Beyond performance enhancement, the ESPD framework makes explicit the fundamental trade-off between resource overhead and operating conditions, replacing cryogenic infrastructure requirements with system-level integration complexity. Although the realization of large-scale ESPD architectures would require further advances in device integration and experimental implementation, the theoretical formulation and parameter studies presented here establish a concrete foundation for the framework and provide concrete guidance for future experimental efforts.
In essence, this work establishes a conceptually complete and theoretically grounded framework for high-performance single-photon detection at room temperature. It demonstrates that intrinsic device-level limitations need not be overcome solely through component replacement, but can instead be effectively compensated by system-level architectural design. As such, the ESPD framework offers a viable pathway toward next-generation photon detection and provides a general methodology for rethinking performance limits in quantum technologies under practical constraints.
Acknowledgements
The author acknowledges the use of Gemini and ChatGPT (accessed December 2025) to, and only to, support language polishing, flow paragraphs, and correct grammar, to improve readability. All AI-generated suggestions were reviewed, revised, and approved by the authors, who take full responsibility for the accuracy and integrity of the work. All insight, results, methods, analysis, and experiments are ordinarily created and owned by the author.
Code Availability